OpenStudy (anonymous):

The region R bounded by the curve 4y=x² and the lines x=4, y=1. Find the volume of the solid formed when R is rotated completely about (b) the y-axis (c) the line y=1(d) the line x=4 the answers should be (b) 18pi (c) 76pi/15 (d)10pi/3

OpenStudy (thomas5267):

Something is wrong in here. \[ x=2\sqrt{y}\\ \begin{align*} \pi\int_0^1(2\sqrt{y})^2dy&=4\pi\int_0^1y\,dy\\ &=2\pi(y^2)|^1_0\\ &=2\pi \end{align*} \] How do you get \(18\pi\)?

OpenStudy (thomas5267):

@Callisto Help!

OpenStudy (anonymous):

it should be x^2=4y

OpenStudy (anonymous):

then looking at the graph the limits of integration are 1 to 4

OpenStudy (anonymous):

To me, for b) you should break it into 2 parts: \(\int_0^2 f(x) dx)) + volume of the cylinder washer from 2 to 4

OpenStudy (anonymous):